The electron of H-atom transits from n=1 to n=4 by absorbing energy. If the energy of n=1 state is -21.8x10^-19 joule, find the energy in n=4 state

Grade:11

2 Answers

Vikas TU
4 years ago
Hiii
We know the formula 
1/Lambda = R (1/nf2-1/ni2)
R = 1.097 * 10^7 meter -1
Lambda = 9.72* 10^-8
E=hc/Lambda
E = 20.450 * 10^-19
So E at n=4 is : -1.4 * 10^-19 joule.
 
Shadow The Hedgehog
4 years ago
Um..sir, can you please give the solution in more details? Because it's a little confusing. Please clarify the solutions by a bit.